OpenStudy (anonymous):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. Part B: write the equation obtained in part A using function notation. Part C: describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

Can any one help me

OpenStudy (mathstudent55):

She paid $285 for 3 days and she paid $510 for 6 days. They want the number of days to be x and the cost to be y. That means we have two ordered pairs: (3, 285) (6, 510) You need to write the equation of the line that passes through those two points.

OpenStudy (mathstudent55):

Do you know how to find the equation of a line when you are given two points on the line?

OpenStudy (anonymous):

Do I have to do slope formula

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Correct. You need to find the slope of the line using the two given points. Can you do that?

OpenStudy (anonymous):

Yea I got 75 @mathstudent55

OpenStudy (anonymous):

As my slope

OpenStudy (mathstudent55):

Correct. Here is the equation of a line given 2 points. The fraction is the slope, so you can put the slope there. \(\Large y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}(x - x_1)\) \(\Large y - y_1 = \dfrac{510 - 285}{6 - 3}(x - x_1)\) \(\Large y - y_1 = 75(x - x_1)\) Now just choose one of the points and replace \((x_1, y_1)\) in the formula with the coordinates of the point you chose.

OpenStudy (anonymous):

What do you mean when you say replace ( X1, Y1) in the formula with coordinates

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

\(\Large y - \color{green}{y_1} = 75(x - \color{red}{x_1})\) Use point \((\color{red}{x_1}, \color{green}{y_1})\) such as \((\color{red}{3}, \color{green}{285} )\) \(\Large y - \color{green}{285} = 75(x - \color{red}{3})\) Now just simplify the expression.

OpenStudy (anonymous):

Ok wait @mathstudent55

OpenStudy (mathstudent55):

ok

OpenStudy (anonymous):

I got y= 75x + 60

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

\(\Large y - 285 = 75(x - 3)\) \(\Large y - 285 = 75x - 225\) \(\Large 75x - y + 60 = 0\) Your equation is correct. Whether your answer is correct or incorrect depends on what your problem interprets "standard form" as.

OpenStudy (mathstudent55):

For part B you need function notation. That simply means solve for y, which you already did. Then replace \(y\) with \(f(x) \).

OpenStudy (anonymous):

Ok and what about C

OpenStudy (mathstudent55):

For part C, think of graphing the equation using the slope-intercept form: y = 75x + 60 Do you know how to graph an equation directly from the slope-intercept form without finding ordered pairs?

OpenStudy (anonymous):

No

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):